OpenStudy (anonymous):

2x+y=1 9x + 3y = -3 The x-coordinate of the point of intersection is

OpenStudy (anonymous):

how do you get rid of the y

OpenStudy (anonymous):

There are two ways to get x. The so-called substitution method: Solve the first equation for y to get y=1-2x, then sub 1-2x into the x in the second equation: 9(1-2x)+3y=-3 Then solve for y, and back sub to get x

OpenStudy (anonymous):

You can multiply the first equation throughout by 3, then subtract the first from the second.

OpenStudy (anonymous):

I would rather multiply the first equation by -3 to get -6x-3y=-3 and add it to the second equation which will eliminate the y: -6x+9x-3y+3y=-3-3 or 3x=-6 or x=-2

OpenStudy (anonymous):

This is called the linear combination method or the elimination method